Exercise Oro Company has two departments with the following production data. It has historically used a single-plant-wide oveneed help with exercise 1 parts 1-3
Answer #1

Ans.

1.

a)

Single Plant-Wide Overhead Rate = Total Manufacturing Overhead Cost / Direct Labor Cost

Total Manufacturing Overhead Cost = $ 120,000 + $ 180,000 = $ 300,000

Direct Labor Cost = $ 25,000 + $ 6,250 = $ 31,250

Single Plant-Wide Overhead Rate = $ 300,000 / $ 31,250 = $ 9.6

b)

Departmental Overhead Rate - Welding = Manufacturing Overhead Cost / Direct Labor hours

Departmental Overhead Rate - Welding = $ 120,000 / 2500 DLhr = $ 48 Per Direct Labor Hour

c)

Departmental Overhead Rate - Fabrication = Manufacturing Overhead Cost / Machine hours

Departmental Overhead Rate - Fabrication = $ 180,000 / 2,000 Machine hours = $ 90 per machine hour

2.

ABC Overhead Rates = Expected Cost / Expected Activity Driver

a ABC Rate - Handling $           9.60 Per part
b ABC Rate - Storing $           1.20 per sq. ft.
c ABC Rate - Inspecting $         48.00 per batch
d ABC Rate - Purchasing $         24.00 per PO
e ABC Rate - Insurance $           0.48 per sq. ft.
f ABC Rate - Cleaning $           0.24 per sq. ft.
g ABC Rate - Utilities $           7.20 per mach. Hour

Cost Assigned = Estimated production data for products * ABC Overhead rate related to it

Standard Deluxe
h Handling costs assigned $    7,200.00 $            4,800.00
i Storage costs assigned $    2,100.00 $               900.00
j Inspecting costs assigned $ 12,000.00 $         12,000.00
k Purchasing costs assigned $ 21,000.00 $            9,000.00
l Insurance costs assigned $    4,800.00 $            1,200.00
m Cleaning costs assigned $    2,400.00 $               600.00
n Utility costs assigned $    3,600.00 $            5,400.00
o Total overhead $ 53,100.00 $         33,900.00
Standard Deluxe
p Overhead Cost per unit $            3.54 $                    6.78

3.

a)

Single plant-wide rate = Total Overhead Cost / Machine Hours

Single plant-wide rate = $ 87,000 / 1,250 Machine hours = $ 69.6 per machine hour

Standard Deluxe
b Total Overhead assigned $ 34,800.00 $         52,200.00
c Overhead Cost per unit $            2.32 $                  10.44